OpenStudy (anonymous):

HELP ME WITH SLOPE!!!!! if i am trying to find slope between two points how do i set it up? i know the formula is y=mx+b but how do i plug in two points for that?

OpenStudy (anonymous):

@phi do you think you can help me?

OpenStudy (anonymous):

@akash_809 here is the other question

OpenStudy (anonymous):

the slope is the tangent of the angle of the line with the horizontal line, try using that with the coordinates and the definition of the tangent of an angle

OpenStudy (aripotta):

the formula for slope is

OpenStudy (anonymous):

im confused i thought the formula was y=mx+b

OpenStudy (aripotta):

that's the formula for a line

OpenStudy (anonymous):

Don't give the answer, let people think first

OpenStudy (anonymous):

If you think about what I said you get that formula

OpenStudy (aripotta):

i didn't give her the answer, i gave her the correct formula

OpenStudy (anonymous):

i dnt even know wat u said @ivanmlerner and he yes he gave me answer because its not like its a math problem

OpenStudy (anonymous):

That is the answer, because she just needs to put the numbers in without thinking

OpenStudy (anonymous):

they dnt make u think in school they just give u the formulas thts wat he did

OpenStudy (anonymous):

I know, and that is wrong

OpenStudy (anonymous):

@AriPotta thanks can u continue to help me please? ignore him ur not just giving me answers because thats not even my main question i needed to know how to set it up to know how to solve my problem

OpenStudy (anonymous):

@ivanmlerner just be quiet becuz if it was wrong they wudnt be teachers and they wudnt be teaching it to us duh!

OpenStudy (aripotta):

ok, what do you need help with?

OpenStudy (anonymous):

i was confuesed on when i use the y=mx+b and when i use the formula u gave me

OpenStudy (aripotta):

in the formula, y=mx+b, m is the slope. find the slope, and you find m.

OpenStudy (anonymous):

i kno if i have two points the how do i find slope? i use the formula u gave me ^?

OpenStudy (phi):

when you have time, this might help http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/algebra--slope

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

so what is b?

OpenStudy (anonymous):

ohk thanks @phi

OpenStudy (aripotta):

the y-intercept